OpenStudy (anonymous):

Two geometric means are inserted between 6 and 384 so the four numbers form a geometric sequence. What are these two numbers? Hint: an = a1(r)n − 1, where a1 is the first term and r is the common ratio

OpenStudy (welshfella):

a1 r^3 = 384 a1 = 6 so r^3 = 384/6 = 64 now you need to find r r = cube root of 64 = ?

OpenStudy (welshfella):

do you follow the above?

OpenStudy (anonymous):

Yes thank you!

OpenStudy (welshfella):

yw